OpenStudy (precal):

The line defined by the equation (1/2)x+3=-2(y-3) is the line tangent to the graph of a function f(x) when x=a. What is the value of f ' (a)?

OpenStudy (precal):

are they looking for just the slope?

OpenStudy (precal):

I solved for y y=(-1/4)x+(3/2) so the slope of the tangent line is -1/4

OpenStudy (precal):

f ' (a)=-1/4 I figured out that a is -6 f '(-6)=-1/4 but I am not sure this is correct or if that is what they are asking for.

OpenStudy (thomas5267):

I think they are just looking for the slope.

hartnn (hartnn):

f'(a) is the slope of tangent to f(x) at x =a

hartnn (hartnn):

yes, f'(a) =-1/4 is correct and thats it.

OpenStudy (precal):

ok thanks...